Square root of x+1 is equal to x-1
Answer by scott8148(6628)

x + 1 = (x - 1)^2

x + 1 = x^2 - 2x + 1

0 = x^2 - 3x = x(x - 3)

x = 0

x - 3 = 0 ___ x = 3